Two 2.00 mC point charges are located on the x-axis. One is at x=1.00 (a) Determine the electric field on the y-axis at y=0.500 m (b) Calculate the electric force on a -3.00C charge placed on the y-axis at y=0.500 m
1
Expert's answer
2020-10-19T13:19:45-0400

A) find the electric field strength of the first charge at the point "y=0.5" m: "E_1=k\frac{q_1}{r^2}=9\times 10^9\times \frac{2\times 10^{-3}}{0.5^2+1}=144\times 10^5" V. According to the parallelogram rule, the sum of two equal stress vectors is

"E_{total}=\sqrt{E_1^2+E_2^2}=""\sqrt{(144\times 10^5 )^2+(144\times 10^5 )^2}\approx 2.04 \times 10^7" V.

B) since the electric field strength is "E=\frac{F}{q}" , where "F" is the Coulomb force, the force acting on the charge "a" is "F=E_{total}\times q_a=2.04 \times 10^7\times 3=6.11\times 10^7" N.